### v1

# A Non-Parametric Sensitivity Analysis for Bounding Bias in Hybrid Control Trials

Abstract
In the digital era it is easier than ever to collect and exploit rich covariate information in trials. Recent work explores how to use this information to integrate external controls, including the use of hybrid control trials (HCTs) where a randomized controlled trial is augmented with external controls. HCTs are of particular interest due to their ability to preserve partial randomization while also improving trial efficiency. However, most HCT estimators rely on an unrealistic assumption: that the external controls are drawn from the same population as the trial subjects (perhaps conditionally on covariates). There has been little formal work to quantify the inevitable bias introduced from a violation of this assumption, slowing the acceptance of HCT designs. To address this, we introduce a non-parametric sensitivity analysis that recognizes that the assumption can be reframed as a ‘no unobserved confounders’ assumption. We leverage omitted variable bias methodologies to estimate the maximum bias introduced from unmeasured covariates, allowing for a critical evaluation of the causal gap that invalidates significant findings. We show that with a relatively weak understanding of the covariate-outcome relationship and of the distinguishability of trial and external subjects, this method reliably bounds bias while also allowing for gains in efficiency. We conclude by discussing considerations for designing and evaluating HCTs, drawing on insights from simulations and theoretical analyses.
1 Introduction
Randomized controlled trials (RCTs) are widely regarded as the gold standard in causal inference because they eliminate confounding, allowing for unbiased estimation [37]. However, RCTs face many limitations, including high costs, ethical constraints, and recruitment difficulties, that make them infeasible or impossible at times [10, 19, 39]. These challenges can compromise research quality through underpowering of trials or inflated type I error rates. Such problems are particularly common in research areas such as rare diseases or vulnerable populations, including pregnant mothers and marginalized populations [6, 23].
Hybrid control trials (HCTs) are an alternative to randomized controlled trials in which an RCT is supplemented with additional control data from external sources (also known as “pooling”). Although not common in practice, the use of HCTs has been proposed as early as 1976 but has since received attention due to their potential to alleviate some of the restrictions mentioned above on RCTs [34, 38, 9]. Regulatory agencies, including the FDA and EMA, have shown a growing interest and commitment to real-world evidence from these trials [6, 34, 9, 1, 4, 5, 3]. The 21st Century Cures Act likewise acknowledges the need to accelerate medical product development in the face of lengthy randomized controlled trials [2]. ‘Borrowing’ information from external sources aligns with these goals and can decrease trial timelines and sample sizes. Furthermore, data availability continues to grow as trials are conducted and pharmaceutical companies participate in data sharing of concurrent control data. Using external controls also offers the opportunity to change the randomization ratio in the RCT portion of the HCT [30]. This allows for more patients to receive the treatment rather than the control, making the trial both more ethical and attractive to prospective participants. With the growth of data, increased support from regulatory agencies, and the existing challenges of traditional RCTs, the time has come to explore data borrowing alternatives such as hybrid control trials.
Multiple data borrowing methods have been proposed with different positions on the bias-variance tradeoff spectrum. At one end of the spectrum, non-HCT methods such as prognostic covariate adjustment train models on previous data to create prognostic scores for trial participants [36, 29]. These do not directly incorporate external data as additional control data but instead use the predictions as an additional covariate to adjust on. Thus, bias is not introduced with the use of the external data. However, since information is borrowed conservatively, the variance reductions are relatively small compared to other methods.
Some HCT methods incorporate external data more directly by ‘pooling’ external and trial control data together. By treating external data as trial data, researchers artificially increase the RCT sample size. Typically, variance decreases and power increases with an increase in sample size. However, the risk of bias is also substantially increased if and when the external population does not truly reflect the trial population. These methods are often referred to as “static” borrowing: gains in efficiency are highest, but robustness against bias is lowest. With static borrowing, researchers always pool the data, assuming that the external data is compatible [42]. The artificial trial sample size increases accordingly to the size of the external sample size, rapidly improving trial power. However, assuming compatibility of the data (which is often unrealistic) introduces estimation bias and type I error inflation.
Even more drastic borrowing methods exist such as single-arm trials where trial participants receive only the treatment, and the control arm consists solely of external data. Note that this is not a form of an HCT as the control arm is not hybrid. Single-arm trials are only acceptable in rare cases since they rely fully upon external data to make comparisons of treatment against control, making them the most susceptible to bias from the use of external data [3, 7].
Besides these “static” borrowing approaches, researchers have also considered “dynamic” approaches [38]. Dynamic borrowing methods fall between the extremes of the bias-variance spectrum, assessing compatibility of control data to adaptively change reliance on external data. These include Bayesian approaches such as power priors and the commensurate prior model [38, 16, 33, 20]. Bayesian power prior methods rely on a weighting power parameter that ranges from 0 (equivalent to an RCT with no external data) to 1 (equivalent to a fully pooled HCT where external and trial data are treated equally). This power parameter is given a prior distribution that then is updated based on the utilized data to optimize weighting of data. These approaches are attractive since they measure differences in the data while still cleverly incorporating external control data. However, large power gains generally cannot be realized without sacrificing type I error control [24, 18].
Frequentist dynamic borrowing includes the test-then-pool approach where data is first assessed for significant differences such as by hypothesis tests [38]. If differences are significant, no pooling occurs; otherwise, pooling proceeds. By testing beforehand, researchers attempt to minimize bias by only pooling when data incompatibility is statistically insignificant. However, if no pooling occurs, researchers solely rely on RCT data with no avenues for gains in efficiency. If pooling occurs, the data are treated as if they are additional trial control units, which results in considerable power gains. However, this approach is criticized for its inability to consistently detect biases, leaving it vulnerable to bias from pooling [27]. Finally, more rigorous frequentist dynamic borrowing methods have been proposed that rely on adaptive lasso penalization to select a subset of suitable external controls [17]. Only external controls whose bias parameters are (or close to) 0 are utilized, resulting in minimized bias from pooling. This method works well compared to others, notably in terms of mean-squared error, especially when differences in control groups are large. However, issues with type I error control arise with finite samples, and gains in power are not as large as other pooling methods due to the conservative nature of selecting suitable controls when differences are already small.
Due to the large potential to increase power, we focus on HCT pooling methods that treat external data similarly to trial data (these include static and frequentist dynamic borrowing methods). The trial-specific average treatment effect (ATE) is the primary parameter of interest in these pooled data experiments, and thus many estimators have been proposed in HCT literature [41, 28, 25]. However, these HCT estimators rely on an unrealistic assumption: mean exchangeability of the controls [28]. A violation of this assumption introduces bias into estimation, potentially inflating type I error rates or decreasing power depending on the direction of the bias. In the context of clinical trials — and particularly of phase III confirmatory trials — these errors are especially problematic. Literature has focused on the development of these estimators, but few, if any, have discussed a formal approach to quantify the bias introduced by an inevitable violation of this assumption, slowing the acceptance of HCTs [41, 1].
1.1 Contributions
To mitigate against this bias, we propose that stringent sensitivity analysis has the potential to maintain the reliability of findings from these trials. We extend research from Chernozhukov et al. [14] where a non-parametric sensitivity analysis leverages efficient debiased machine learning and produces easily interpretable bounds on omitted variable bias in observational studies. In theory, the mean exchangeability assumption is met when all relevant covariates are controlled for confounding. It follows that a violation of the mean exchangeability assumption is a result of omitted variable bias. We modify this sensitivity analysis in the context of HCTs to reliably bound bias given plausible increases in the explanatory power from the omitted variables on the “Riesz representer” and outcome regression of the trial-specific ATE without making any additional assumptions on the data generating process [14]. Results of the sensitivity analysis can then be compared against the strength of evidence from HCTs to determine whether conclusions are still significant even after accounting for the potential bias.
The aim of this research is to expand upon the current literature exploring the practicality of HCTs in the clinical setting. By proposing a reliable assumption-free sensitivity analysis for the pooled estimator, we add credibility to results from HCTs, unlocking the benefits they have to offer. We believe that increasing acceptance of these trial designs will accelerate the development of new therapies, particularly benefiting under-funded or underrepresented areas of research.
Outline.
In Section 2 we begin by defining necessary notation and preliminaries for the hybrid control trial setting. We also explore a doubly robust efficient estimator for the trial-specific average treatment effect [28]. In Section 3 we validate that it is possible to extend the omitted variable bias sensitivity analysis to the HCT setting. Through identification of the Riesz representer and outcome regression of the trial-specific ATE, we derive an estimator for the bias bound $B$. In Section 4, we use simulations to demonstrate that the risk of bias and error is minimized with the sensitivity analysis and that trial efficiency can be improved even when bias is present. We do this by showing (i.) that the true bias is reliably bounded by $B$, (ii.) that the sensitivity analysis corrects inflated type I error rates in HCTs, and (iii.) that power can be increased when using the HCT estimator combined with the sensitivity analysis. In Section 5, we summarize key findings and explore guidelines for designing and evaluating HCTs based on these results as well as the potential impacts of this research.
2 Notation and Preliminaries
2.1 Notation
Consider a hypothetical hybrid control trial, consisting of a randomized controlled trial and an external control arm. Let the variable $D$ denote trial participation with $D=1$ indicating that a subject was a participant in the randomized controlled trial and $D=0$ indicating that a subject was not a participant in the randomized controlled trial. Equivalently, $D=0$ implies the participant is part of the external control arm. Let the variable $Y$ represent the outcome of interest in the study and $X$ represent the measured covariate(s) in the study. Let $A$ represent the binary treatment of interest with $A=1$ indicating that a subject received the treatment of interest and $A=0$ indicating that a subject did not receive the treatment. Note that since the external data only consist of control data, a subject with $D=0$ implies $A=0$. The observed data follow structure $O=(D,A,X,Y)$. Furthermore, we assume that the data $O_{1},...,O_{n}$ are i.i.d. distributed from a true data distributing distribution $\operatorname*{\mathbb{I\kern-1.49994ptP}}_{0}$.
Variable Definition
$D$ Trial participation indicator: $D=1$ if in RCT, $D=0$ if external control.
$A$ Treatment indicator: $A=1$ if treated, $A=0$ otherwise.
$X$ Measured covariate(s).
$Y$ Observed outcome, with $Y=A\,Y_{1}+(1-A)\,Y_{0}$.
[…]
$Y_{1}$ Counterfactual outcome under treatment ($a=1$).
$Y_{0}$ Counterfactual outcome under control ($a=0$).
$Z$ Unobserved confounders (s.t. $Y_{a}\perp\!\!\perp A\mid X,Z$).
$O^{*}$ Full (latent) data structure $(D,A,X,Z,Y_{1},Y_{0})$.
Table 1: Random variables and their definitions.
We also consider the unobserved full dataset. Let $Y_{1}$ and $Y_{0}$ denote the counterfactual outcomes where $Y_{a}$ represents the outcome that would be observed under treatment $a\in\{0,1\}$. We define the observed outcome $Y$ under treatment $A$ with the counterfactual outcomes $Y_{a}$ such that $Y=A\cdot Y_{1}+(1-A)\cdot Y_{0}$ (others call this the “causal consistency” assumption, often bundled with the IID assumption as “SUTVA”). Assume that there is a set of unobserved random variables Z such that $Y_{a}\perp\kern-9.0pt\perp A|X,Z$. These represent our “unobserved confounders” without which we cannot guarantee mean exchangeability. The full data structure follows as $O^{*}=(D,A,X,Z,Y_{1},Y_{0})$, which is i.i.d. distributed from some distribution $\operatorname*{\mathbb{I\kern-1.49994ptP}}_{0}^{*}$. We can think of $\operatorname*{\mathbb{I\kern-1.49994ptP}}_{0}$ as marginal relative to $\operatorname*{\mathbb{I\kern-1.49994ptP}}_{0}^{*}$ over $Z$ and with only $Y$ observed rather than $Y_{1}$ and $Y_{0}$.
Let $N_{d}$ with $d\in\{0,1\}$ denote the sample size of the trial and external datasets for $d=1$ and $d=0$ respectively. Since we are assuming $D$ is i.i.d. distributed from $\operatorname*{\mathbb{I\kern-1.49994ptP}}_{0}$, the external sample size and RCT sample sizes are not fixed. The total sample size of the hybrid control trial $n$ is the sum of $N_{0}$ and $N_{1}$.
Let $\mu(A,X)=E[Y|X,D=1,A=a]$ for $a\in\text{support}(A)$ define the trial-specific outcome regression. Let $p(X)=P(A=1|X,D=1)$ define the trial-specific treatment propensity score — note that the non-trial-specific treatment propensity score is always 0 since all external data consists of control units. Let $q=P(D=1)$ be the marginal probability of trial participation with $q=E[\frac{N_{1}}{n}]$. Let $\pi(X)=P(D=1|X)$ define the selection propensity score; this serves a similar role as the treatment propensity score but instead estimates the conditional probability of RCT trial participation given covariates $X$. Finally, let $r(X)=\frac{V[Y|X,D=1,A=0]}{V[Y|X,D=0,A=0]}$ represent the variance ratio of the outcomes between the concurrent trial controls and external controls.
Parameter Definition
$\mu(A,X)$ $E\bigl{[}Y\mid X,D=1,A=a\bigr{]}$, the trial‐specific outcome regression.
$p(X)$ $P\bigl{(}A=1\mid X,D=1\bigr{)}$, the trial‐specific treatment propensity score.
$q$ $P(D=1)\;=\;E\!\Bigl{[}\tfrac{N_{1}}{n}\Bigr{]}$, the marginal probability of trial participation.
$\pi(X)$ $P\bigl{(}D=1\mid X\bigr{)}$, the selection propensity score.
$r(X)$ $\frac{V\bigl{[}Y\mid X,D=1,A=0\bigr{]}}{V\bigl{[}Y\mid X,D=0,A=0\bigr{]}}$, the conditional outcome variance ratio of concurrent vs. external controls.
Table 2: Population parameters and their definitions.
2.2 The Trial-Specific Average Treatment Effect
In this paper, we focus on the trial-specific average treatment effect. This measures the causal effect of the treatment $A$ on outcomes $Y$ for the population of people who were eligible to participate in the randomized controlled trial. Although other estimands of interest exist for hybrid control trials, we focus on the trial-specific average treatment effect due to its prevalence in hybrid control trial causal inference literature [28, 41, 40]. This parameter is also of key importance in the pharmaceutical and regulatory context where RCT trial recruitment is based on a carefully selected population, so inference would likewise need to be focused on this specific population.
The causal trial-specific average treatment effect is:
$\Psi^{*}=E[Y_{1}-Y_{0}|D=1]$
This estimand parallels the typical average treatment effect that would be estimated in a standard RCT. Two key assumptions are needed for identification [^1] Some authors consider causal consistency to be an identification assumption as well. Although we do “assume” causal consistency, we think of it more as a structural link between the causal and observable distributions rather than as a restriction on the set of causal data generating processes. The latter are what we prefer to call “identification assumptions” but this is ultimately a matter of perspective. of the trial-specific average treatment effect in our HCT setting:
  1. 1.
    Strong Ignorability:
    1. (i)
      Conditional Independence: $Y_{a}\perp\kern-9.0pt\perp A|X,D=1$ for all $a\in A$
    2. (ii)
      Positivity: $P(A=1|X=x,D=1)>0$ for all $x\in X$
  2. 2.
    Mean Exchangeability of $Y_{0}$: $E[Y_{0}|X=x,D=1]=E[Y_{0}|X=x,D=0]$ for all $x\in X$
Strong ignorability is a common assumption to identify the average treatment effect when conducting inference in observational studies. That condition is met by design for the purposes of our paper since the trial is randomized. On the other hand, the mean exchangeability assumption is not addressed by randomization, leaving a gap between the observable estimand that can be estimated and the causal estimand of scientific interest. This gap is precisely the bias induced by pooling in HCTs.
Using the above assumptions, the following observable statistical estimand identifies the causal trial-specific ATE (see A.1 for proof):
$\Psi=E[E[Y|A=1,X,D=1]-E[Y|A=0,X]]$
An important benefit of the mean exchangeabilty assumption is that although we are interested in the trial-specific average treatment effect (thus the situation where $D=1$), the second term in the estimand above is not conditioned on $D$. Indeed, the mean exchangeability assumption ensures that $E[Y|A=0,X]=E[Y|A=0,X,D=1]=E[Y|A=0,X,D=0]$ in the context of strong ignorability. Due to this, the conditional expectation of the control subjects may be “evaluated” over the population of all possible HCT participants (not just the trial population). In other words, we can use the external control subjects to help estimate the trial-specific ATE even though they are not exactly from the population of interest.
2.2.1 Efficient Estimation
We rely on an efficient estimator $\hat{\Psi}$ for the trial-specific average treatment effect proposed by Li et al. [28]. Although others exist, we focus on this estimator because it is an efficient estimator. This property aligns with our goal to alleviate trial constraints since efficient estimators permit smaller sample sizes while preserving standard methods of inference and robust behavior across data-generating processes [25]. This estimator is also doubly robust, making it less prone to model misspecification bias.
The efficient influence function for this estimand (as derived by Li et al. [28]) is as follows:
$\phi(Y,A,X,D)$ $=\frac{1}{q}\left[D(\mu(1,X)-\mu(0,X)-\Psi)+\left(\frac{DA}{{p}(X)}-{W}(A,X,D)\right)(Y-\mu(A,X))\right]$
${W}(A,X,D)$ $=\frac{D(1-A)+(1-D){r}(X)}{{\pi}(X)(1-{p}(X))+(1-{\pi}(X)){r}(X)}\pi(X)$
The derivation of this influence function takes into account that $P(A=0|D=0)=1$, i.e. that the external data contains only controls. Explicitly incorporating this knowledge actually increases the efficiency of the analysis (i.e. is expected to give smaller confidence intervals and p-values).
Solving an empirical score equation based on this influence function yields the following efficient estimator for the trial-specific average treatment effect (with $E_{n}$ estimating the population expectation with a sample average):
$\hat{\Psi}$ $=\frac{1}{\hat{q}}E_{n}\left[D(\hat{\mu}(1,X)-\hat{\mu}(0,X))+\left(\frac{DA}{\hat{p}(X)}-\hat{W}(A,X,D)\right)(Y-\hat{\mu}(A,X))\right]$
$\hat{W}(A,X,D)$ $=\frac{D(1-A)+(1-D)\hat{r}(X)}{\hat{\pi}(X)(1-\hat{p}(X))+(1-\hat{\pi}(X))\hat{r}(X)}\hat{\pi}(X)$
Given that we are selecting $q$ by design of the HCT experiment, the estimation of $\hat{\Psi}$ depends primarily on the estimation of functional components $\mu(A,X)$, $p(X)$, and $\pi(X)$ as well as the variance ratio $r(X)$. In the formula for the efficient estimator, $\hat{\mu}(A,X)$, $\hat{p}(X)$, $\hat{\pi}(X)$, and $\hat{r}(X)$ are machine learning estimates of their corresponding population objects, converging at assumed $L^{2}$ rates.
Under general conditions, $\hat{\Psi}$ is asymptotically normal with large-sample variance of $\sqrt{n}\hat{\Psi}$ given by $E[\phi(O)^{2}]$, which can be approximated with a plug-in estimator using the appropriate estimated regressions. This facilitates the construction of confidence intervals and p-values.
When one assumes mean exchangeability and has a binarized outcome $Y$, $r(X)$ reduces to 1. In this paper, we will restrict the outcome to be binary so that $r(X)$ does not require estimation and the results can be presented more economically. However, the results should still extend well to other cases (e.g. continuous or even censored outcomes). For example, the value of $r(X)$ only affects efficiency and not unbiased estimation, and $r(X)$ can be estimated well in a scenario where a continuous outcome is of interest as shown in Li et al. [28].
The estimator given above is a typical efficient estimating equations (EEE, or “AIPW-style”) estimator. Although we use EEE here for accessibility and consistency with the double machine learning literature, other (asymptotically equivalent) approaches for efficient estimation such as targeted maximum likelihood estimation (TMLE) can also be used [25].
3 Methods
The issue at hand is that this estimator, like others, can only target the observable estimand which relies on the strict mean exchangeability assumption in order to get any benefit from the external control subjects. Although there are approaches such as the test-then-pool method to detect violations of this assumption, they are typically simple and cannot recognize every violation [42]. Further, preliminary tests may identify a violation of the mean exchangeability assumption, but it’s also possible that the violations only lead to small amounts of bias. In these cases, the external data may have been useful, providing greater benefits than costs, but not used due to the black-and-white nature of significance tests. More complicated methods of dynamic borrowing include Bayesian methods, but simulation studies have shown that power gains are limited when strict type-one error rates are enforced [24]. Due to the inevitability and uncertain quantification of this bias, hybrid control trials are typically discouraged by major regulatory bodies [1].
Considering these issues, we shift focus away from if bias is present to how much bias is present when this assumption is violated. We argue that the benefits of data borrowing in HCTs can lead to overall increased trial feasibility even if the external data are slightly biased away from the trial data. Therefore, there is great interest in quantifying this bias and bounding type-one error inflation so that researchers can safely reap the benefits of hybrid control trials, including smaller sample sizes, shorter timelines, and reduced budgets. This motivates the following analyses.
3.1 Omitted Variable Bias
Any violation of the mean exchangeability assumption can be seen as a direct result of omitted variable bias. In the instance that $E[Y_{0}|X,D=1]\neq E[Y_{0}|X,D=0]$, we should imagine that mechanistically there is some unobserved confounder $Z$ that induces the difference in the outcome’s dependence on the observed covariates.
As an extreme and unrealistic example, consider a hypothetical hybrid control trial where a randomized controlled trial is conducted today and the external controls are collected from a previous trial. Suppose that this external data was collected far in the past, such as a century ago. Even if they share information on covariates $X$ such as age, gender, weight, etc., it is no surprise if the average outcomes in a particular covariate stratum differ between the trial controls and external controls, i.e. $E[Y_{0}|X,D=1]\neq E[Y_{0}|X,D=0]$. One can deduce that there may be confounding simply due to the stark differences between the modern time and the time when the external data were collected. However, as relevant information that captures these time- and place-dependent differences is collected — such as sanitary conditions, occupations, quality of diet, exposure to pollutants, among many others — the conditional expected outcomes should begin to converge. If all relevant confounders, consisting of the sets $X$ and $Z$ (originally unmeasured), are somehow able to be measured, the conditional expected outcomes will become equivalent between our control groups. We can call this full set of confounders $\tilde{X}$.
This idea generalizes to any hybrid control trial and recontextualizes the issue to be an omitted variable bias problem. In particular, $\Psi$ no longer identifies $\Psi^{*}$ since it does not include the relevant covariates. Instead, $\tilde{\Psi}$, denoting the statistical estimand that conditions on $\tilde{X}$, identifies the true causal parameter $\Psi^{*}$. This discrepancy introduces the bias $|\Psi^{*}-\Psi|=|\tilde{\Psi}-\Psi|$. By bounding this bias, we can estimate how far $\Psi$ is from $\tilde{\Psi}$ when unobserved confounding is present. In the rest of this paper, we also use the tilde to denote long regressions (e.g. $\tilde{\mu}(A,\tilde{X})=E[Y|\tilde{X},A=a,D=1]$) in which the full set of confounders $\tilde{X}$ is used instead of the short regressions in which only the observed covariates $X$ are used.
In what follows, our goal will be to find an upper-bound $|\Psi^{*}-\Psi|\leq B$, quantifying the maximum bias that could be introduced if the mean exchangeability assumption were violated. We will then derive an estimator $\hat{B}$ for $B$, the bias bound.
3.2 The Bias Bound $B$
We leverage prior omitted variable bias methodologies for developing sensitivity analyses based on the Riesz representation of the causal estimand [14]. Prior sensitivity analyses have been developed for a variety of estimands in observational studies such as the average treatment effect and weighted average incremental effects. In particular, this method works for a class estimands that can be written as a linear functional of the conditional expectation function of the outcome, also known as the outcome regression. These estimands can be written in the form: $\Psi(P_{0})=E[m(\mu)(O)]$ where $\mu$ is the outcome regression based on $P_{0}$ and $m$ is some linear functional.
In the case of the trial-specific ATE, we can use conditioning arguments to write the estimand as $E[m(\mu)(O)]$ where $m(\mu)$ returns the function $O\mapsto q^{-1}D(\mu(1,X)-\mu(0,X))$ (see A.2 for proof).
3.2.1 Riesz Representation
The Riesz representation theorem states that within a Hilbert space any continuous linear function that maps to a real number can be written as the inner product of the argument to the function and some other element in the Hilbert space [35, 26]. Using this theorem, we can then rewrite estimands of the form given above as $E[m(\mu)(O)]=E[\mu(O)\alpha(O)]$ where $\alpha$ is defined as the Riesz representer.
As an example, we can consider the well-known estimand for the average treatment effect of a binary treatment: $\Psi(P_{0})=E[E[Y|A=1,X]-E[Y|A=0,X]]$. The ATE can be defined as a linear functional of the outcome regression with $m(\mu)(O)=\mu(1,X)-\mu(0,X)$. Standard conditioning arguments show that the Riesz representer for the ATE is $\frac{A}{P(A=1|X)}-\frac{1-A}{P(A=0|X)}$, the usual inverse propensity weights. In this way, the Riesz representation motivates the use of such weighted estimators [26].
Chernozhukov et al. [14] uses this representation to define a bias bound $B$ for any Riesz-representable estimand, dependent on the residuals between the long and short formulations of the outcome regression and Riesz representer. Since the short regressions are projections of the long regressions onto a smaller space of functions depending only on $X$, these short regressions are orthogonal to the additional regression components living outside of this smaller subspace. Relying on the orthogonality of $\mu$ and $\alpha$ on these residuals, it is shown that the bias (in terms of the difference of the long and short estimands) is bounded: $\Psi-\tilde{\Psi}=E[(\mu(O)-\tilde{\mu}(O))(\alpha(O)-\tilde{\alpha}(O))]\leq\|\mu-\tilde{\mu}\|\|\alpha-\tilde{\alpha}\|$. That is: the maximum bias is the covariance of the approximation errors of the short representations, which is upper bounded by the product of the $L^{2}$ errors. The implication of this result is that there is a sharp bias bound that doesn’t depend on the full complexity of the (“long”) data-generating distribution, only on how well the short regressions approximate the long regressions in mean-squared error.
We are able to extend these findings to the hybrid control trial scenario in which we use the trial-specific average treatment effect, which is also a partially linear functional of the conditional expectation function of the outcome. The following fits the trial-specific ATE into the methodology proposed by Chernozhukov et al. [14].
3.2.2 Deriving the Riesz Representer
First, we identify the Riesz representer using the Riesz representation theorem, showing that $\Psi=E[E[Y|A=1,X,D=1]-E[Y|A=0,X]]$ is equal to the inner product of the input outcome regression $\mu(A,X)=E[Y|A=a,X,D=1]$ (also equivalent to $E[Y|A=a,X,D=0]$ by the mean exchangeability identification assumption when A=0) and some Reisz Representer $\alpha(A,X,D)$. In A.3, it is shown that the Riesz representer of the trial-specific average treatment effect is
$\alpha(A,X,D)=\frac{1}{q}\left(\frac{DA}{p(X)}-\frac{D(1-A)+(1-D){r}(X)}{{\pi}(X)(1-{p}(X))+(1-{\pi}(X)){r}(X)}\right)\pi(X).$
3.2.3 Measuring Unobserved Confounding
The methodology proposed by Chernozhukov et al. [14] relies on the outcome regression and the Riesz representer, which have been identified above, and gains in explanatory power from latent, or unobserved, confounders. We can formally define $C^{2}_{Y}$ and $C^{2}_{D}$ as the gain in information from including unobserved confounders $Z$ in the outcome regression and Riesz representer in terms of the non-parametric $R^{2}$, letting the tilde denote the long regression in which the unobserved confounders are included in estimation of the functional components.
$C^{2}_{D}$ $=\frac{E[\tilde{\alpha}(A,X,D)^{2}]-E[\alpha(A,X,D)^{2}]}{E[\alpha(A,X,D)^{2}]}$
$C^{2}_{Y}$ $=\frac{E[(\tilde{\mu}(A,X)-\mu(A,X))^{2}]}{E[(Y-\mu(A,X)^{2}]}$
3.2.4 Calculating the Bias Bound $B$
As in Chernozhukov et al. [14], we reformulate the squared-bias as
$|\Psi-\tilde{\Psi}|^{2}$ $=\rho^{2}C^{2}_{Y}C^{2}_{D}S^{2}$
$\ S$ $=\sqrt{E[Y-\mu(A,X)]^{2}E[\alpha(A,X,D)]^{2}}$
$\rho$ $=Cor(\mu-\tilde{\mu},\alpha-\tilde{\alpha})$
where $\rho$ measures the degree to which confounding in $m$ is related to confounding in $\alpha$.
The quantity $S$ depends only the observable (“short”) distribution, whereas the other parameters depend on the long distribution. Therefore only $S$ can be estimated from the data. $\rho$, $C_{D}$, and $C_{Y}$ must be assumed or bounded on the basis of domain knowledge. There are, however, sensible ways of setting these parameters. It is possible to pick a good value for $\rho$ — as well as $C^{2}_{D}$ and $C^{2}_{Y}$ — empirically through benchmarking methods using the observed covariates. For example, given a dataset with multiple observed covariates, one can select an observed covariate to be treated as unobserved and calculate estimates for $m$, $\tilde{m}$, $\alpha$, and $\tilde{\alpha}$ to then use to estimate $\rho$, $C^{2}_{D}$, and $C^{2}_{Y}$ through a “leave-one-out” approach [32]. The decision of which variable to treat as unobserved depends on multiple factors such as the nature of confounding we might expect to be present as well as the potential magnitude of confounding. However, there is no free lunch: any method for choosing these parameters incurs some risk and therefore requires a good understanding of the problem domain. In practice these parameters should also be varied over plausible ranges — see Section 3.4 on critical causal gaps below.
In our paper, we proceed with the most extreme estimate of $|\rho|=1$ for a conservative estimate of $B$. As a result, our bias bound is
$B=C_{Y}C_{D}S.$
Under adversarial confounding when $\rho^{2}=1$, the bias bound is sharp with $B=C_{Y}C_{D}S$. In our simulations, we proceed with the assumption that $C_{D}$ and $C_{Y}$ have been well-chosen a priori (e.g. using one of the methods described above) since we want to separately evaluate the performance of our method from a user’s knowledge or assumptions. Certainly, our method will appear to fail if these parameters are poorly chosen.
3.2.5 EEE Estimator of $B$
In order to estimate $B$, we must first estimate $S$, which depends on our Riesz representer and outcome regressions. For the outcome regression $\mu(A,X)$, we estimate separate regressions for each treatment. To estimate $\mu(1,X)$, we use only the data corresponding to those who were part of the original RCT, regressing $Y$ onto $A$ and $X$ for only $D=1$. However, due to the mean exchangeability assumption we can estimate $\mu(0,X)$ with both the RCT and external data arms (thus where we gain efficiency in the estimation of the trial-specific ATE), regressing $Y$ onto $A$ and $X$ regardless of $D$. Although pooling the data to estimate $\mu(0,X)$ when mean exchangeability does not hold introduces bias, this is precisely the bias that we are measuring in this paper. If we can discern the magnitude of the bias, we can evaluate whether the risk of bias from pooling outweighs the benefit of increased efficiency from pooling.
To estimate $\alpha(A,X,D)$, we use a plug-in estimator based on the functional components outlined in Section 2. We first use machine learning to estimate $\pi(X)$ and $p(X)$ by respectively regressing $D$ onto $X$ for all participants and $A$ onto $X$ only for those who participated in the RCT. We can then use these to estimate $\alpha(A,X,D)$ for every participant in the HCT, plugging in values into the equation for $\alpha$.
Each of the regression tasks we describe can be accomplished with generic nonparametric or machine learning methods, e.g. random forests, boosting, neural networks, kernel ridge regression, elastic net, etc. When we have estimated $\alpha$ and $\mu$, we can then plug these in to estimate $S$ such that $\hat{S}=\sqrt{E_{n}[Y-\hat{\mu}(A,X)]^{2}E_{n}[\hat{\alpha}(X,A,D)]^{2}}$. This can then be used in the estimation of the bias bound, plugging in $\hat{S}$ with plausible hypothetical values for $C_{Y}$ and $C_{D}$.
3.2.6 Inference on $B$
Since the bias bound requires estimation of the term $S$, it is itself a random quantity with some uncertainty that should be quantified. Inference on $B$ (via $S$) can be conducted using the delta method to combine estimates of the standard errors of $\nu^{2}=E[\alpha(X,A,D)]^{2}$ and $\sigma^{2}=E[Y-\mu(A,X)]^{2}$. The influence functions for $\nu^{2}$ and $\sigma^{2}$ are $\phi_{\sigma^{2}}=(Y-\mu(A,X))^{2}-\sigma^{2}$ and $\phi_{\nu^{2}}=\alpha(X,A,D)^{2}-\nu^{2}$. See Chernozhukov et al. [14] for additional details. The influence function of $B$ is then:
$\varphi_{B}=\frac{|\rho|}{2}\frac{C_{Y}C_{D}}{S}(\sigma^{2}\phi_{\nu^{2}}+\nu^{2}\phi_{\sigma^{2}}).$
Given the influence function, the variance of $\sqrt{n}B$ is $E[\varphi_{B}(O)^{2}]$. We estimate variance with a simple plug-in estimator, using estimated versions of each of the required population quantities (e.g. $\hat{\psi}_{\nu^{2}}=\hat{\alpha}(X,A,D)^{2}-\hat{\nu}^{2}$).
3.3 Using the Bias Bound
Presuming that we had a true upper bound $B\geq b$ and the true value of the observable trial-specific ATE $\Psi$, we could define a range $[\Psi_{-},\Psi_{+}]$ (calculated using $\Psi_{\pm}=\Psi\pm B$) which would be certain to include the true, causal, trial-specific ATE. Our task is therefore to produce estimates and conduct inference for the estimands $\Psi_{\pm}$ using our efficient estimates of $\Psi$ and $B$ which are described above.
3.3.1 Inference on $[\Psi_{-},\Psi_{+}]$
Since the target has changed from estimating $\Psi$ to $\Psi_{\pm}$, inference likewise changes. The influence functions of the bounds $\Psi_{\pm}$ can be derived using the delta method, relying on the influence functions of $\Psi$ and $B$, as follows:
$\varphi_{\pm}=\phi\pm\varphi_{B}$
where $\phi$ is the influence function for $\Psi$ defined in section 2. The variance of the upper and lower bounds, $\sigma^{2}_{+}$ and $\sigma^{2}_{-}$, can thus be found taking the variance of $\varphi_{\pm}$.
Instead of constructing a two-sided confidence interval around the estimates of the bounds (such as one would do for inference on $\hat{\Psi}$), we follow Imbens and Manski [22] to construct one-sided confidence bounds on $\Psi_{+}$ and $\Psi_{-}$ that maintain uniform asymptotic coverage across $B$.
$CI_{\eta}=\left[\hat{\Psi}_{-}-\frac{c_{\eta}\hat{\sigma}_{-}}{\sqrt{n}},\hat{\Psi}_{+}+\frac{c_{\eta}\hat{\sigma}_{+}}{\sqrt{n}}\right]$
with $c_{\eta}$ such that
$\Phi\left(c_{\eta}+\frac{\sqrt{n}(\hat{\Psi}_{+}-\hat{\Psi}_{-})}{\max\{\hat{\sigma}_{-},\hat{\sigma}_{+}\}}\right)-\Phi(-c_{\eta})=1-\eta$
[…]
Given a critical value of $\eta$, this confidence region that uses both confidence bounds guarantees coverage of $\Psi$ to asymptotically approach $1-\eta$ uniformly in $B$. This is important when there is very little unmeasured confounding. Another benefit of this procedure is that whenever $B\approx 0$, the confidence region becomes equivalent to that of a naive HCT analysis with a standard two-sided pointwise confidence interval. More details and justification can be found in Imbens and Manski [22].
3.4 Identifying the Critical Causal Gaps
In reality, the true values of $C^{2}_{Y}$ and $C^{2}_{D}$ will always be unknown since they depend on the unobserved confounders $Z$. Based on different combinations of $C^{2}_{Y}$ and $C^{2}_{D}$, it is possible to identify values of $C^{2}_{Y}$ and $C^{2}_{D}$ where any significant results from the HCT analysis are invalidated by the potential magnitude of bias present. These are values that result in the confidence bounds of $\hat{\Psi}_{+,-}$ crossing the null value (in many cases this is 0, indicating that there is no trial-specific average treatment effect). A robustness value can also be calculated as a summary of robustness against bias [14]. The robustness value measures the minimum strength of confounding required for both $C^{2}_{D}$ and $C^{2}_{Y}$ to negate the significant findings of the HCT. For example, if the robustness value is 0.1, we know that any combinations of $C^{2}_{Y}$ and $C^{2}_{D}$ where both are greater than 0.1 will result in insignificant results. Thus, when the robustness value is relatively low, it indicates there is a large risk of bias from using the external control data.
We can again extend the work of Chernozhukov et al. [14] to create contour plots to display these combinations of $C^{2}_{Y}$ and $C^{2}_{D}$, such as the following:
Figure 1: A contour plot demonstrating robustness of a hypothetical HCT. The x-axis shows increases in explanatory power from Z in the Riesz Representer, and the y-axis shows increases in explanatory power from Z in the outcome regression. The unadjusted (naive) HCT analysis lies at the origin and the dotted red line establishes the threshold where $B$ negates significant findings. Different combinations in $C^{2}_{D}$ and $C^{2}_{Y}$ can also be specified, as indicated by the red diamond. Here, we use the robustness values as the scenario, demonstrating one point at which results become insignificant. The lower confidence bound of the trial-specific ATE is shown in parentheses.
Given a similar plot, researchers should be able to identify plausible increases in $R^{2}$ for the outcome regression and the Riesz representer through professional knowledge of the topic or other methodologies for omitted variable bias. Some of these methodologies include algorithms such as the “leave-one-out” approach introduced earlier and others explored in section 5. Conclusions can then be made based on the robustness of evidence from the HCT at the researcher’s discretion. Although it may be impractical to assume that researchers can pinpoint the exact amount of unobserved confounding, it is plausible that researchers can assess robustness against bias due to the flexible nature of the contour plots. If the data are shown to not be robust against bias, it may be best to proceed with a pure RCT analysis. Particularly, if the plausible increases in $R^{2}$ far exceed the critical causal gap values, the risk of introducing bias exceeds the potential benefits of incorporating external data.
3.5 Summary
In summary, the steps one would take to execute this analysis are as follows:
  1. 1.
    Gather external control data with measurements on the same outcome variable and covariate variables as the RCT data. Through some preliminary analysis (examples in Section 5), check that the covariate distribution of the two sets of control data are not significantly dissimilar.
  2. 2.
    Use double debiased machine learning to estimate $S$ and $\Psi$ as described above.
  3. 3.
    Pick $\rho=1$ (for conservative estimates) and for plausible values or ranges of $C^{2}_{Y}$ and $C^{2}_{D}$:
    1. (a)
      Calculate the bounds $\Psi_{\pm}$ using the estimates of $\Psi$, $S$, and posited values of $C^{2}_{Y}$ and $C^{2}_{D}$.
    2. (b)
      Estimate one-sided confidence bounds using the adjustment for when $B\approx 0$ and standard error of the bounds found by combining the variances of $\Psi$ and $B$.
  4. 4.
    Optional: Create contour plots to visualize results and/or calculate robustness values for flexible interpretation.
4 Simulation Study
In the following section, we use simulations to verify the performance of the sensitivity analysis as well as to explore the practicality of aiming for improved efficiency even with the wider confidence bounds resulting from the sensitivity analysis. In particular, it is shown that the true bias is correctly and consistently bounded by the bias bound $B$ in varying scenarios, type-one error inflation is corrected by the sensitivity analysis, and power can be improved even after accounting for potential omitted variable biases.
4.1 Data Generating Process
The following data generating process (DGP) was used:
[…]
$Y^{*}|A,X,Z$ $=9.5+\beta A-0.45X+0.75Z+\mathcal{N}(0,1)$
$Y$ =$=\begin{cases}1\text{ when }Y^{*}\geq 10\\ 0\text{ when }Y^{*}<10\end{cases}$
Note that the DGP is not meant to be realistic, but rather be relatively simple so that the operating characteristics of the sensitivity can be understood under easily-manipulated scenarios. In this scenario, Y is a binarized outcome, indicating that some measurement of interest $Y^{*}$ was greater than or equal to a cutoff value of 10. A binary outcome was chosen to proceed with the simplified formulation of the Riesz representer $\alpha$. Several values in the simulation are variable: $q,\zeta,\beta$, allowing manipulation for different scenarios. As defined earlier, $q$ is the probability of trial participation, which equates to $E[\frac{N_{1}}{n}]$. Throughout all simulations, $q$ is varied from 0.1 to 0.9 with step size 0.1. $\zeta$ is the probability of some binary unobserved confounder Z for external control participants. Bias can be manipulated by leveraging the distribution of the unobserved confounder $Z$ and the size of the external control arm $N_{0}$. $\beta$ is the effect of treatment on the measurement of $Y^{*}$. By changing this value, we are able to assess scenarios in which there is a true treatment effect and in which there is a true null treatment effect.
4.2 Simulation Setup
For the following simulations, R 4.4.2 was used with seed 509.
[…]
For all algorithms, we rely on double debiased machine learning using the ranger algorithm [13]. Double debiased machine learning was used for the estimation of $\Psi$, $\Psi_{ate}$, and $S$ (the component in the bias bound not dependent on the parameters $\rho$, $C^{2}_{D}$, and $C^{2}_{Y}$). Random forests were chosen to estimate the functional components of these statistical estimands due to their ability to perform quickly and well without much tuning.
A key assumption used in the simulations is that the true values of $C^{2}_{Y}$ and $C^{2}_{D}$ are known. Using the DGP, we can closely estimate these values by generating hypothetical trials with 1,000,000 participants. The non-parametric $R^{2}$ can then be compared between scenarios in which only $X$ is used for estimation and in which the full set of relevant confounders $\tilde{X}$ are used to find the true values of $C^{2}_{Y}$ and $C^{2}_{D}$ using the formulation in Section 3. While this is not possible in real HCT trials, the aim of the simulations is to show that the sensitivity analysis is reliable, given we have some sufficient understanding of the covariate-outcome relationship.
While the simulations that follow use specific values of RCT and external arm sample sizes, the sample sizes of these groups in each iteration are ultimately determined by the randomness of the DGP and $q$. For example, if we want to explore a scenario in which 100 RCT participants are desired, the total HCT sample size $n$ is first found by dividing 100 by $q$. Then the DGP randomly assigns the $n$ participants to either the RCT or external arm. Thus, the RCT sample size fluctuates loosely around the desired sample size of 100. Note that the rest of the paper refers to these desired RCT and external sample sizes with the understanding that each iteration will roughly fall near these desired sample sizes since these are not actually fixed in the DGP.
4.3 Verifying the Bias Bound
In Figure 2, the terms “strong bias”, “medium bias”, and “weak bias” are used. These indicate how different the distribution of the unmeasured confounder $Z$ is between the trial participants and the external control participants. The RCT sample size is set to roughly 100 with varying amounts of external data, determined by $q$. The total sample size can then be calculated by dividing the input RCT sample size by each value of $q$. The sensitivity analysis was applied to the data to obtain estimates of $B$ and their corresponding 95% confidence intervals for each simulated scenario. To calculate the true bias $b$, 10,000,000 hypothetical participants were generated with the DGP to calculate $b=|E[Y_{1}-Y_{0}|D=1]-(E[E[Y|A=1,X,D=1]-E[Y|A=0,X]])|$.
Scenarios:
Variable Weak Bias Medium Bias Strong Bias
$\zeta$ 0.4 0.6 0.9
$\beta$ 0.75 0.75 0.75
Table 3: Values used in the simulation for Figure 2. The strength of bias is varied by manipulating the underlying distribution of $Z$ in the external control group in comparison with the underlying distribution of $Z$ in the RCT group
4.3.1 Bias can be correctly estimated with any amount of external data but increases as external data size grows
[…]
The estimation of the bias bound $B$ in the sensitivity analysis consistently correctly bounds the true bias $b$ with $B>b$ in every simulated scenario, regardless of the ratio of external data to trial data and the strength of bias arising from $Z$.
As expected, the size of the true bias $b$ increases both as the difference in the distributions of $Z$ increases and the size of the external control arm increases. The increase in bias as the external control arm increases can be explained by the increase in amount of biased data, increasing the influence of the bias. The bias bound $B$ follows a similar shape in response to the trajectories of $b$. Interestingly, the standard error of the bias bound $B$ also increases as the size of the external control arm increases. Since the standard error of the bias bound is partially a function of the expectation of $\alpha$, it is driven by the $\frac{1}{q}$ term, inflating as $q$ becomes closer to 0 (equivalently as $n_{0}$ increases relative to $n_{1}$). Given this, one should be hesitant on using an overwhelming amount of external controls in comparison to the RCT sample size and even more so if the true effect size is expected to be relatively small.
4.4 Type I Error Control
[…]
Scenarios:
Variable No Bias Bias
$\zeta$ 0.3 0.4
$\beta$ 0 0
Table 4: Values used in the simulation for Figure 3. The presence of bias is varied by manipulating the underlying distribution of $Z$ in the external control group in comparison with the underlying distribution of $Z$ in the RCT group
4.4.1 Inflated type I error rates are corrected with the sensitivity analysis.
Figure 3: Type I error rates are compared across three trial setups: pure RCT, HCT, and HCT combined with the sensitivity analysis (shown in blue, red, and green respectively). Multiple scenarios are tested, varying the presence of bias and external sample sizes. From left to right: external sample size of 100, external sample size of 500, and external sample size of 1000. From top to bottom: bias present, no bias present. RCT sample sizes are varied with fixed external sample sizes.
Focusing first on the scenario where there is no bias, the sensitivity analysis returns the same results as the naive HCT. Since there is no bias, $B=0$ and calculations for confidence bounds return the same intervals as calculations for confidence intervals for naive HCTs. Here we can check that the coverage for the HCT trial-specific estimator is similar to that for the RCT ATE estimator without need to worry about the influence of bias. Despite no bias, the type I error rate of HCT analysis is slightly inflated in smaller RCT sample sizes of 50 when there are larger external sample sizes of 500 and 1000. However, we see that the errors of the HCT estimator approach the error rates of the RCT estimator as RCT sample size increases (as quick as $N_{1}=100$). We believe this reflects the nature of the second-order remainders of both since they are different between the trial-specific ATE and the RCT ATE. In Appendix B.1 we provide a theoretical explanation for why type I errors of the trial-specific estimator may inflate when the ratio of external to trial data is large. However, as reflected in Figure 3, these differences quickly disappear.
[…]
4.5 Gains in Efficiency
[…]
Scenarios:
Variable No Bias Favorable Bias Unfavorable Bias
$\zeta$ 0.3 0.4 0.2
$\beta$ 0.6 0.6 0.6
Table 5: Values used in the simulation for Figure 4. The favorability of bias is varied by manipulating the underlying distribution of $Z$ in the external control group in comparison with the underlying distribution of $Z$ in the RCT group
4.5.1 Hybrid controls can be used to mitigate underpowering of small RCTs even in the presence of bias
Figure 4: Power curves are compared across analysis based on pure RCT, naive HCT analysis, and HCT analysis combined with the sensitivity analysis (shown in solid blue, solid lines, and dotted lines respectively). External sample size is fixed in panels while RCT sample size is varied. From left to right: external sample size of 100, external sample size of 500, and external sample size of 1000. Bias scenarios are varied based on favorable bias, unfavorable bias, and no bias (shown in red, purple, and green).
Simulations show that even in the presence of bias, HCTs can improve trial efficiency compared to underpowered RCTs. Across all simulated scenarios, power is always improved when an RCT is supplemented with external data. Note that this may not always be the case, especially if the external data is more different from the RCT control data than simulated, in which case it is possible that any gains in efficiency will be negated by large biases. However, in scenarios where the bias bound $B$ is estimated to be relatively large to the point that there is little benefit to using external controls, the contour plots and critical causal gaps described in Section 3 would likely return results that advise researchers to proceed with a pure RCT analysis. As a result, it is unlikely that there would be many scenarios in which the sensitivity analysis results in worsened underpowering of trials (as researchers would simply use the RCT analysis, resulting in the same power). We also see that the amount of external data is also important as the dotted lines pull further away from their respective solid lines as the external sample size grows, aligning from findings from figure 2 where larger external sample sizes result in an increase in $b$. However, in our simulated scenarios it seems as though this is usually offset by the gains in efficiency from the external data.
[…]
4.6 Conclusion
These results suggest that the use of external controls combined with the sensitivity analysis is a practical solution in relieving traditional RCT trial constraints, many of which stem from sample size requirements, even if the unrealistic mean exchangeability assumption is not met. These results also address and remedy the traditional complaints against naive HCTs where biases can greatly influence estimation and inference, notably inflating type I error rates.
5 Discussion
The findings in the Section 4 imply that there needs to be careful deliberation in selecting the ideal ratio of RCT data against the external control data. From Figure 3, we see finite sample size issues with type I error inflation when data are heavily imbalanced, even when there is no bias present. This finding stems from the form of the second-order remainder of the trial-specific ATE as described in Appendix B.1 where it takes longer for the asymptotics to ‘kick in’ compared to the RCT ATE when $\frac{1}{\hat{q}}$ is small. We believe this to be an important finding, as the form of the remainder has not been explored in prior trial-specific ATE literature, despite its relevance to understanding the asymptotic behavior of estimators under certain conditions, such as varying values of $q$. However, regardless of the external sample sizes in the simulations, these issues seem to disappear once the RCT sample size reaches 100. As a result, we recommend being wary of using unbalanced ratios when RCT sample size is especially small. However, as seen in Figure 4, efficiency can be safely improved through the combined use of external data and sensitivity analysis even when using smaller external sample sizes.
When RCT sample size is sufficiently large (100$\approx$ 100≈ 100), the possibility of using larger external sample sizes arises. The question of how large depends on several factors. First, we must consider the expected effect size of treatment, and then we must consider how different the external control arm may be from the RCT control arm. However, this second consideration should not be a major concern given proper selection of the external control arm. In fact, it would be improper to proceed with HCT analysis if the external control arm were expected to be significantly different than the RCT control arm. Despite this, it is worth considering the size of the inevitable bias from confounding since it can be seen in Figures 2 and 4 that any difference is amplified when external sample size is increased. The expected effect size is then important here; if the expected effect size is relatively small, amplified bias threatens correctly identifying a treatment effect. In this case, we recommend using a smaller external control arm. However, in the case that the expected effect size is relatively large (such as in the simulations), we see in Figure 4 that we can safely increase power by increasing the external sample size. The greatest power gains were achieved in the simulations when external sample size was 500 or 1000. While the gains in power are large from 100 to 500, the gains in power are relatively smaller from 500 to 1000. Consequently, it may be best to still limit the ratio of data away from extreme imbalances since the potential consequences outweigh the marginal benefits of relying heavily on external data.
These considerations, however, all rely on the assumption made in the simulations that the true values of $C^{2}_{Y}$ and $C^{2}_{D}$ are known. In reality, the size of confounding due to omitted variable bias is never truly known. As mentioned above, omitted variable bias and sensitivity literature has explored different benchmarking algorithms to determine the relative plausible sizes of unobserved confounding [8, 14, 15, 21, 11]. In particular, benchmarking methods are popular where the observed covariates are used to approximate the influence of unobserved confounders. It has been argued that given strategic selection of measured covariate variables, the unobserved covariates likely are equally or less explanatory of outcome and treatment assignment [8]. With this assumption, one can use observed covariates to estimate bounds on the relationships of any unobserved variables with outcome and treatment. McClean et al. [32] summarizes these methods where notable strategies include the “leave-one-out” approach, addressed earlier in the methods section, as well as more complicated methods such as “leave-some-out” and “leave-every-out” to also rely on subsets of observed covariates. In all, these algorithmic approaches hold promise as benchmarking methods and can be used in combination with this sensitivity analysis to estimate plausible ranges of values for $C^{2}_{Y}$ and $C^{2}_{D}$ in order to judge robustness against omitted variable bias. While benchmarking itself provides measures of confounding due to unobserved confounders, critics have pointed out that relying solely on benchmarking may lead to erroneous conclusions [15, 32]. The benefit of using these benchmarks in conjunction with the sensitivity analysis outlined here is that the bias is properly identified, leading to safer conclusions on the causal effect of treatment [15].
Finally, the scope of this paper focuses on binary outcomes for simplicity. However, we expect these findings to generalize easily to the continuous outcome scenario. The main motivation behind focusing on binary outcomes was the $r(X)$ term, the ratio of variance in the outcome between the RCT control data and external control data, in the Li estimator for the trial-specific average treatment effect. To reduce the number of values to be estimated, we binarized the outcome such that $r(X)$ would equal 1 under mean exchangeability. For continuous outcomes where $r(X)$ would need to be estimated, the original paper provides insight on empirical methods [28]. Further, we believe that the results would translate to continuous outcomes since the variance ratio is expected to only affect efficiency, not bias. Therefore, the idea of increasing trial efficiency while protecting against bias with the sensitivity analysis should be able to be extended to continuous outcomes after some verification with simulations.
For estimation, we relied on an efficient estimating equations approach with debiased machine learning. However, so long as the framework in the paper is followed, other estimation techniques can also be used. For example, methods such as targeted maximum likelihood estimation (TMLE) may be applied after solving for the canonical gradient of the bias bound [14, 25]. We however relied on an EEE approach for an intuitively understandable approach by prospective researchers who may implement the sensitivity analysis. Further, we relied on the Li estimator for the trial-specific ATE due to its efficiency, but the sensitivity analysis can be applied for other estimators of the trial-specific ATE so long as the Riesz representer and outcome regression can be properly identified.
For selection of the external control population, it is important that the external control subjects are as similar as possible to the RCT control subjects to maintain the integrity of results. Several guidelines exist for selection of external controls, such as Pocock’s selection criteria [34]. In particular, the external control subjects must have:
  1. 1.
    Received the same treatment as the randomized controls
  2. 2.
    Been selected from a recent clinical study with comparable eligibility criteria
  3. 3.
    Been evaluated in the same method as the randomized controls
  4. 4.
    Had comparable distributions of important covariates as the randomized controls
  5. 5.
    Been selected from a study conducted by the same organization with overlap in clinical investigators
  6. 6.
    No indications of significant differences between the external and randomized controls.
These strict criteria attempt to ensure that the external controls are as similar as possible to the randomized control subjects. However, it is argued in some literature that these can be relaxed in extreme cases where data is limited or there is a pressing need to conduct research [30]. For example, there has been movement towards also including electronic health record (EHR) data as a source of external control data [12]. The implication of adding EHR data as an external data source is large, vastly enriching the availability of recent, relevant health data. Regardless of a relaxation of these criteria, data selection should still mirror these criteria as closely as practically possible to balance minimizing biased decision making and minimizing the patient burden.
Other attempts to minimize bias from external control selection include a priori approaches such as a test-then-pool approach [42]. In this approach, a simple hypothesis test is conducted on the outcome prior to the decision to pool the external and randomized controls. Given an insignificant finding, researchers pool the data together and analyze the data as an HCT. Given a significant finding, the hypothesis test raises a red flag, resulting in researchers analyzing the RCT separately from the external data. Li et al. [28] also suggest a test for the mean exchangeability assumption in which a parametric outcome model is fit with interactions between $D$ and $X$. A variety of other tests exist in which this mean exchangeability assumption can be tested in advance of pooling such as the non-parametric omnibus test of equality [31] However, these approaches are all limited in that they are only asymptotically guaranteed to detect discrepancies in the data. Since there will always be a level of bias between the control groups, these tests will always choose to not pool the data asymptotically. However, these are applied to small samples in the HCT setting where the availability of data (from the RCT and/or external data source) is constrained. As a result, these will only identify any significantly large discrepancies in the data. The sensitivity analysis is thus a logical complement to such tests (and pooling) to detect the smaller inevitable differences in the data that lead to bias in naive HCT analysis.
The ability to safely use external control data in hybrid control trials has multiple implications. HCTs have been proposed as early as 1976, but this trial design has typically been discouraged by major regulatory bodies [1]. Despite openness to the use of external data, the reliance on the mean exchangeability assumption, which practically is always violated to some degree, has hindered the wide acceptance of HCTs. HCT literature has predominantly focused on either creating new, more efficient estimators or relaxing the mean exchangeability assumption. Bayesian methodologies have been proposed, in which one can dynamically adjust how much information to borrow from the external data based on compatibility of the controls, but gains in efficiency are minimal when enforcing strict type I error control [24]. However, HCT literature has not explored what one can do when the mean exchangeability assumption is not assumed to be met. This paper fills in this gap in the research, quantifying this inevitable bias through a sensitivity analysis. We hope that filling in this gap in the research and making findings more credible will lead to the wider acceptance of HCTs.
The benefits of the HCT trial design are numerous. Of course, there is reduced recruitment, leading to shorter timelines and budgets, but the benefits go beyond these factors. Ethically, the HCT trial design is attractive as it minimizes patient burden [30]. A study relying on solely an RCT requires more patients to be randomized, resulting in a larger number of patients missing out on receiving a treatment. In addition, with increased control information, it is possible to change the randomization ratio such that more RCT participants receive the treatment than the control. This not only is more ethical but can also be more attractive to prospective participants, encouraging trial participation. Another benefit of adding credibility to HCTs is that it can decrease reliance on single-arm study designs in which all control data come from external sources. Single-arm trials are also used in small sample settings, but are even more dependent on the mean exchangeability assumption, resulting in an even higher risk of bias [38]. While there are cases in which even an HCT may be infeasible, the cases in which an HCT is feasible should provide a safer alternative trial design than the single-arm trial [30].
Considering its extreme nature, HCTs are only recommended in situations where there is a pressing need to borrow information from outside of an RCT. In particular, this includes rare disease research, oncology trials, and traditionally excluded populations from research. Due to the nature of these research areas, conducting a full-scale RCT is usually impractical, hindering the development of medicines and treatments for these populations. The findings of this paper align with the aims of the 21st Century Cures Act “to accelerate the discovery, development, and delivery of 21st century cures” [2]. With added credibility to the findings from HCTs, regulatory bodies will hopefully allow for more HCT trials, greatly increasing the volume of research that can be dedicated to these traditionally difficult areas in medical research.
Appendix A Proofs
A.1 Identification Proof
$E[Y_{1}|D=1]$ $=E[E[Y_{1}|X,D=1]]$
$=E[E[Y|A=1,X,D=1]]$ (1)
$E[Y_{0}|D=1]$ $=E[E[Y_{0}|X,D=1]]$
$=E[E[Y|A=0,X,D=1]]$ (1)
$=E[E[Y|A=0,X]]$ (2)
$E[Y_{1}-Y_{0}|D=1]$ $=E[Y_{1}|D=1]-E[Y_{0}|D=1]$
$=E[E[Y|A=1,X,D=1]]-E[E[Y|A=0,X]]$
$=E[E[Y|A=1,X,D=1]-E[Y|A=0,X]]$
Assumptions: (1) Ignorability, (2) Mean Exchangeability of Controls
A.2 Linear Functional of Trial-Specific ATE
$\Psi$ $=E[E[Y|A=1,X,D=1]-E[Y|A=0,X]]$
$=E[E[E[Y|A=1,X,D=1]-E[Y|A=0,X]|D=1]]$
$=E[\frac{D}{q}(E[Y|A=1,X,D=1]-E[Y|A=0,X]]$
$=E[\frac{D}{q}(\mu(1,X)-\mu(0,X)]$
$=E[m(\mu)(O)]$
Where $m(\mu)(O)$ is defined as $\frac{D}{q}(\mu(1,X)-\mu(0,X))$ in the main text.
A.3 Riesz Representer of Trial-Specific ATE
$qE[E[Y|A=1,X,D=1]]$ $=E[\pi(X)E[Y|A=1,X,D=1]]$
$=E\left[\pi(X)\frac{D}{\pi(X)}E[Y|A=1,X,D]\right]$
$=E\left[\pi(X)\frac{D}{\pi(X)}\frac{A}{p(X)}E[Y|A,X,D]\right]$ (1)
$=E\left[\frac{DA}{p(X)}E[Y|A,X,D]\right]$
$q[E[E[Y|A=0,X]]]$ $=E[\pi(X)E[Y|A=0,X]]$
$=E\left[\pi(X)E[Y|A=0,X,D=1]\frac{\pi(X)(1-p(X))}{\pi(X)(1-p(X))+(1-\pi(X))r(X)}\right.$
$\quad\left.+\pi(X)E[Y|A=0,X,D=0]\frac{(1-\pi(X))r(X)}{\pi(X)(1-p(X))+(1-\pi(X))r(X)}\right]$ (2, 3)
$=E\left[\pi(X)\frac{D(1-A)}{\pi(X)(1-p(X))}E[Y|A,X,D]\frac{\pi(X)(1-p(X))}{\pi(X)(1-p(X))+(1-\pi(X))r(X)}\right.$
$\quad\left.+\pi(X)\frac{(1-D)}{1-\pi(X)}E[Y|A,X,D]\frac{(1-\pi(X))r(X)}{\pi(X)(1-p(X))+(1-\pi(X))r(X)}\right]$ (1)
$=E[W(A,X,D)E[Y|A,X,D]]$
$E[E[Y|A=1,X,D=1]-E[Y|A=0,X]]$ $=E[E[Y|A=1,X,D=1]]-E[E[Y|A=0,X]]$
$=\frac{1}{q}E\left[\frac{DA}{p(X)}E[Y|A,X,D]\right]-\frac{1}{q}E[W(A,X,D)E[Y|A,X,D]]$
$=\frac{1}{q}E\left[\underbrace{\left(\frac{DA}{p(X)}-W(A,X,D)\right)}_{\text{Riesz Representer}}E[Y|A,X,D]\right]$
Assumptions: (1) Positivity, (2) Total Expectation, (3) Mean Exchangeability of Controls
Note that since we only use mean exchangeability for the control units, we also use our knowledge that $P(A=1|X,D=0)=0$ and $P(A=0|X,D=0)=1$.
Appendix B Second-Order Remainders
B.1 Second-Order Remainder of Trial-Specific ATE
$R$ $=\hat{\Psi}-\Psi+E[\hat{\phi}(Y,A,X,D)]$
$=\hat{\Psi}-\Psi+E\left[\hat{\alpha}(A,X,D)(Y-\hat{\mu}(A,X)+m(\hat{\mu})(O)-\frac{D}{\hat{q}}\hat{\Psi}\right]$
$=\hat{\Psi}-\Psi+E\left[\frac{1}{\hat{q}}\hat{\beta}(A,X,D)(Y-\hat{\mu}(A,X)+\frac{1}{\hat{q}}n(\hat{\mu})(O)-\frac{D}{\hat{q}}\hat{\Psi}\right]$
$=\hat{\Psi}-\frac{1}{\hat{q}}\hat{\Psi}ED+\frac{1}{\hat{q}}E\hat{\beta}(A,X,D)Y-\frac{1}{\hat{q}}E\hat{\beta}(A,X,D)\hat{\mu}(A,X)+\frac{1}{\hat{q}}En(\hat{\mu})(O)-\Psi$
$=\hat{\Psi}-\frac{q}{\hat{q}}\hat{\Psi}+\frac{1}{\hat{q}}E\hat{\beta}(A,X,D)\mu(A,X)-\frac{1}{\hat{q}}E\hat{\beta}(A,X,D)\hat{\mu}(A,X)+\frac{1}{\hat{q}}E\beta(A,X,D)\hat{\mu}(A,X)-\Psi$
$=\left(1-\frac{q}{\hat{q}}\right)\hat{\Psi}+\frac{1}{\hat{q}}E[-\hat{\beta}(A,X,D)\hat{\mu}(A,X)+\hat{\beta}(A,X,D)\mu(A,X)+\beta(A,X,D)\hat{\mu}(A,X)]-\Psi$
$=\left(1-\frac{q}{\hat{q}}\right)\hat{\Psi}+\frac{1}{\hat{q}}E[-\hat{\beta}(A,X,D)\hat{\mu}(A,X)+\hat{\beta}(A,X,D)\mu(A,X)+\beta(A,X,D)\hat{\mu}(A,X)$
$-\beta(A,X,D)\mu(A,X)]+\frac{1}{\hat{q}}E\beta(A,X,D)\mu(A,X)-\Psi$
$=\left(1-\frac{q}{\hat{q}}\right)\hat{\Psi}-\frac{1}{\hat{q}}E(\hat{\beta}(A,X,D)-\beta(A,X,D))(\hat{\mu}(A,X)-\mu(A,X))+\frac{1}{\hat{q}}E\beta(A,X,D)\mu(A,X)-\Psi$
$=\left(1-\frac{q}{\hat{q}}\right)\hat{\Psi}-\frac{1}{\hat{q}}E(\hat{\beta}(A,X,D)-\beta(A,X,D))(\hat{\mu}(A,X)-\mu(A,X))+\frac{1}{\hat{q}}E\beta(A,X,D)\mu(A,X)$
$-\frac{1}{q}E\beta(A,X,D)\mu(A,X)$
$=\frac{1}{\hat{q}}(\hat{q}-q)\hat{\Psi}-\frac{1}{\hat{q}}E(\hat{\beta}(A,X,D)-\beta(A,X,D))(\hat{\mu}(A,X)-\mu(A,X))+(\frac{1}{\hat{q}}-\frac{1}{q})E\beta(A,X,D)\mu(A,X)$
$=\frac{1}{\hat{q}}(\hat{q}-q)\hat{\Psi}-\frac{1}{\hat{q}}E(\hat{\beta}(A,X,D)-\beta(A,X,D))(\hat{\mu}(A,X)-\mu(A,X))+\frac{1}{q}(\frac{q}{\hat{q}}-1)E\beta(A,X,D)\mu(A,X)$
$=\frac{1}{\hat{q}}(\hat{q}-q)\hat{\Psi}-\frac{1}{\hat{q}}E(\hat{\beta}(A,X,D)-\beta(A,X,D))(\hat{\mu}(A,X)-\mu(A,X))-\frac{1}{\hat{q}}(\hat{q}-q)\Psi$
$=\frac{1}{\hat{q}}(\hat{q}-q)(\hat{\Psi}-\Psi)-\frac{1}{\hat{q}}E(\hat{\beta}(A,X,D)-\beta(A,X,D))(\hat{\mu}(A,X)-\mu(A,X))$
Letting $n(\mu)(O)=D(\mu(1,X)-\mu(0,X))=qm(\mu)(O)$ and $\Psi=E(\frac{1}{q}\beta(A,X,D)\mu(A,X))=E(\alpha(A,X,D)\mu(A,X))$ under the Riesz Representation Theorem. Note that here we have manipulated two terms — $m(\mu)(O)$ to $\frac{1}{q}n(\mu)(O)$ and $\alpha(A,X,D)$ to $\frac{1}{q}\beta$ — to pull out the $\frac{1}{q}$ to explicitly highlight how the second-order remainder also depends on estimation of $q$.
Given the consistency of $\hat{\Psi}$ and the central limit theorem applied to $\hat{q}$, the first term is $o_{P}(n^{-1/2})$ as long as $q\neq 0$. The second term in the remainder is a typical “doubly robust” term: $\frac{1}{\hat{q}}E(\hat{\beta}(A,X,D)-\beta(A,X,D))(\hat{\mu}(A,X)-\mu(A,X))$. Typically for second-order convergence of the double robust term, we require that $\|\hat{\mu}-\mu\|\|\hat{\beta}-\beta\|$ is $o_{P}(n^{-\frac{1}{2}})$ [35]. In our setting we must also require $q$ bounded away from 0. This is not an issue.
However, if $q$ is small we should expect a relatively larger impact of this term for fixed $n$, i.e. the asymptotics will not “kick in” until a relatively larger sample size. This behavior can be seen in figure 3 when external sample size is large but RCT sample size is small, leading to inflated type I error rates.
Appendix C Type I Error Rate under Two Repetition Cross Fitting
Figure 5: Type I error rates are compared across three trial setups: pure RCT, HCT, and HCT combined with the sensitivity analysis (shown in blue, red, and green respectively). Multiple scenarios are tested, varying the presence of bias and external sample sizes. From left to right: external sample size of 100, external sample size of 500, and external sample size of 1000. From top to bottom: bias present, no bias present. RCT sample sizes are varied with fixed external sample sizes.